# RALAD: Bridging the Real-to-Sim Domain Gap in Autonomous Driving with Retrieval-Augmented Learning

### v1

@@ Abstract
In the pursuit of robust autonomous driving systems, models trained on real-world datasets often struggle to adapt to new environments, particularly when confronted with corner cases such as extreme weather conditions. Collecting these corner cases in the real world is non-trivial, which necessitates the use of simulators for validation. However, the high computational cost and the domain gap in data distribution have hindered the seamless transition between real and simulated driving scenarios. To tackle this challenge, we propose Retrieval-Augmented Learning for Autonomous Driving (RALAD), a novel framework designed to bridge the real-to-sim gap at a low cost. RALAD features three primary designs, including (1) domain adaptation via an enhanced Optimal Transport (OT) method that accounts for both individual and grouped image distances, (2) a simple and unified framework that can be applied to various models, and (3) efficient fine-tuning techniques that freeze the computationally expensive layers while maintaining robustness. Experimental results demonstrate that RALAD compensates for the performance degradation in simulated environments while maintaining accuracy in real-world scenarios across three different models. Taking Cross View as an example, the mIOU and mAP metrics in real-world scenarios remain stable before and after RALAD fine-tuning, while in simulated environments, the mIOU and mAP metrics are improved by 10.30% and 12.29%, respectively. Moreover, the re-training cost of our approach is reduced by approximately 88.1%. Our code is available at https://github.com/JiachengZuo/RALAD.git.
@@ I INTRODUCTION
Figure 1: The performance of baseline models in CARLA: All models exhibit varying degrees of performance degradation in 3D object detection.
@@
However, autonomous driving systems inevitably encounter corner cases, such as extreme weather conditions, unexpected pedestrian behavior, and rare road scenarios, which challenge their perception and decision-making capabilities [14, 15]. Given the high safety standards required in vehicle operation to prevent life-threatening accidents, it is imperative that autonomous systems rigorously address and test these corner cases to ensure robust performance [14, 15]. Models trained on real-world datasets predominantly encounter regular driving scenarios, making it challenging to cover and address the rare and complex corner cases [21, 16, 19]. To replicate and test corner cases, simulators are often used, as real-world data is difficult to collect [21]. Models trained on real-world datasets cannot be directly applied in simulators because of the gap between real and simulated environments, as shown in Figure 1 and Table I, where we tested three models (Monolayout [5], Cross-View [6], and DcNet [7]) in CARLA, resulting in significant decreases in both performance and accuracy. With the advent of large models, the cost and time required for model training have significantly increased [4]. Therefore, finding cost-effective methods to improve model performance in simulators has become a major challenge.
We found that in the field of computer vision (CV), there are already methods that utilize nearest neighbor search and feature fusion for domain adaptation to address cross-domain dataset problems [16, 18]. Our problem can also be regarded as a cross-domain problem between real datasets and simulated datasets. Given the similarity between their problem and ours, we attempt to apply this method to autonomous driving. However, applying this method to autonomous driving needs to face three major challenges. Firstly, the significant differences between images in other fields and autonomous driving images hinder the direct use of this method in this field, as autonomous driving needs to consider both the direct distance between individual images and the group distance between image sets. Secondly, unlike imaging that completely uses real data, simulated images are completely different from real images. We do not know the best granularity for calculating similarity in simulated data. Finally, it is crucial to meet the need for low-cost training while ensuring the robustness and reliability of the model.
To address these challenges, we propose Retrieval-Augmented Learning for Autonomous Driving (RALAD). To handle the gap between real and simulated images in autonomous driving, we introduce the optimal transport method that considers both individual and group distances. For the unknown optimal granularity in simulated data, we adjust the optimal transport from object-level to pixel-level retrieval. Finally, we adopt a fine-tuning approach by freezing computationally expensive neural network layers and re-training only selected layers, which reduces computational costs while maintaining model robustness and reliability in low-cost training environments. We selected three 3D object detection models for our experiments, as 3D object detection is one of the most critical tasks in autonomous driving. Conducting experiments in this domain provides the strongest justification for evaluating the effectiveness of our approach.
@@
  • We introduce RALAD, a framework that addresses the real-to-sim gap in autonomous driving and provides pixel-level OT capabilities.
  • We apply RALAD to three models, achieving significant performance improvements.
  • We establish a mapping between real and simulated environments and conduct extensive experiments to validate the approach.
Figure 2: The left panel shows the pixel-level retrieval enhancement process: feature extraction by an Encoder, OT algorithm for transport cost calculation, nearest neighbor retrieval, feature fusion, and fine-tuning. The top right details the OT computation process. The bottom right explains fine-tuning with fused features, freezing all but the Decoder layers.
@@ II-A Retrieval-Augmented Learning
Retrieval-Augmented Learning (RAL) is an approach that enhances learning models, it integrates retrieval mechanisms to leverage existing data representations, thereby improving performance and efficiency. For example, Yottixel [22] employs a mix of supervised and unsupervised methods, including segmentation, clustering, and deep networks, to analyze image patches and employ distance metrics for efficient search and retrieval. SISH [23] utilizes a tree structure for rapid WSI search and an uncertainty-based ranking for retrieval, reducing storage and labeling by building on preprocessed mosaics without pixel or ROI labels, using self-supervised learning indices and pre-trained embeddings. HHOT [24] introduces optimal transport (OT) as a metric for comparing whole slide images (WSIs) or across WSI datasets, theoretically underpinning the application of OT for steering the retrieval and assembly of datasets. In RAM-MIL [18], the attention weight serves as a measure of probability density, signifying the ”mass” being transferred. By quantifying this, the method computes the conversion cost across various data domains. It then employs this distribution for nearest-neighbor retrieval, seamlessly integrating features from distinct domains to address out-of-domain challenges.
@@ II-B Gap Between Real And Sim
In the field of autonomous driving, there exists the issue of the Gap during the application process from simulation to reality [25], where discrepancies in lighting, textures, vehicle dynamics, and agent behaviors between virtual and real environments complicate the direct application of simulation results. To address this, researchers have developed two primary approaches: sim2real knowledge transfer and the use of digital twins (DTs) [17, 28]. In knowledge transfer learning for autonomous driving, the RG problem is compounded by uneven environmental sampling and complex physical parameters [26]. To overcome this, researchers have developed strategies such as curriculum learning, meta-learning, knowledge distillation, robust reinforcement learning, domain randomization, and transfer learning [27]. Domain randomization, in particular, helps align simulation parameters with real-world variability, facilitating the transfer of learned strategies to real-world applications. Conversely, digital twin technology creates virtual models of real-world entities or systems. A case in point is the development of the SynFog dataset [28], which uses an end-to-end simulation process to produce photo-realistic synthetic fog data, enhancing learning-based algorithm research and facilitating the model’s transition from synthetic to real data. Nonetheless, despite their promise, these methods confront the issue of high computational expenses, particularly in complex, dynamic real-world settings.
@@
3D object detection is crucial for autonomous driving as it enables vehicles to accurately perceive and understand their surroundings, which is essential for safe navigation and decision-making. Conventionally, this has been achieved with the help of LiDAR sensors, which, although precise, are prohibitively expensive and computationally demanding. To address these limitations, the field has seen a significant advancement with the application of deep learning techniques that leverage monocular cameras for 3D detection. Specifically, the development of bird’s-eye view (BEV) representations from monocular images has emerged as a promising and more cost-effective alternative. Techniques such as MonoLayout [5] and Cross View [6] demonstrate the potential of using these BEV representations to perform 3D object detection. Cross View, in particular, has introduced a cross-view transformation module and a context-aware discriminator to enhance results, achieving cutting-edge performance in vehicle occupancy estimation. Nonetheless, issues such as class imbalance and low computational efficiency remain. The Dual-Cycled Cross-View Transformer network (DcNet) [7] has been proposed to tackle these challenges by integrating focal loss and optimizing multi-class learning, setting new standards of performance in the field of 3D object detection for autonomous driving.
III METHOD
This section provides a detailed explanation of the RALAD framework, as shown in Figure 2. We introduce Pixel-Level Retrieval-Augmented Learning based on Optimal Transport into autonomous driving, using Cross-view as an example to illustrate the RALAD process.
@@ III-A Problem Formulation
The gap between real-world and simulation (real2sim) primarily arises from the cross-domain challenges between real and simulated data in autonomous driving. To address this, we consider two datasets: the Real dataset $D_{r}$ and the Sim dataset $D_{s}$. The Real dataset is defined as $D_{r}=\left\{X_{n},Y_{n}\right\}_{n=1}^{N_{r}}$, where $N_{r}$ is the number of images, and the Sim dataset is $D_{s}=\left\{\tilde{X_{m}},\tilde{Y_{m}}\right\}_{m=1}^{{N_{s}}}$, with ${N_{s}}$ being the number of simulated images. To bridge the gap between these two datasets, the first step is to establish a mapping relationship between them. We extract features from real and simulated images using an encoder function g(·), resulting in $H_{k}$ = g($x_{k}$) for real images and $\tilde{H}_{k}$ = g($\tilde{x}_{k}$) for simulated images, where $x_{k}$ and $\tilde{x}_{k}$ are individual images from $X$ and $\tilde{X}$, respectively. The features $H_{k}$ and $\tilde{H}_{k}$ are composed of pixel-wise feature vectors, with $H_{k}=\left\{h_{i}\right\}_{i=1}^{w*h}$ and $\tilde{H}_{k}=\left\{\tilde{h}_{i}\right\}_{i=1}^{w*h}$, where $h_{i}$ and $\tilde{h_{i}}$ represent the feature vectors of pixel $i$ in the real and simulated images, respectively. Our ultimate goal is to identify the similarity between $H_{k}$ and $\tilde{H}_{k}$, thereby enabling cross-domain retrieval between the two datasets.
III-B Pixel-Level Retrieval-Augmented Learning based on Optimal Transport in Real2Sim
To address the mapping relationship between features from different domains, we introduce Retrieval-Augmented Learning based on Optimal Transport. However, previous RAL approaches were primarily object-level and focused on real-world images. Given the cross-domain nature of our real and simulated environments, simple object-level approaches are insufficient for effectively establishing relationships between real and simulated images. Therefore, pixel-level computations are required to ensure fine-grained matching between real and virtual images. This precise matching aids in identifying and preserving detailed information, thereby establishing a more accurate correspondence between the two domains.
@@ III-B Pixel-Level Retrieval-Augmented Learning based on Optimal Transport in Real2Sim
Furthermore, previous RAM based on OT was applied to simple classification tasks, whereas our autonomous driving BEV perception task is far more complex. To address this challenge, we propose to treat every pixel as a sample in the OT computation, while each image is treated as a probability distribution. We assign uniform weights to all instances and leave the non-uniform OT calculation to future exploration. The purpose of the OT algorithm is to calculate the distance between two features $H_{n}$ (real-world) and $\tilde{H_{m}}$ (simulation) for subsequent retrieval of the nearest feature. The formula is as follows:
$\operatorname{d_{OT}}\left(H_{n},\tilde{H_{m}}\right)=\sum_{i=1}^{w\cdot h}c\left(h_{i},\tilde{h_{i}}\right)T_{ii}+\beta\cdot\sum_{i}T_{ii}\log T_{ii}$ (1)
@@
In this equation, T denotes the transport plan matrix where each element $T_{ii}$ specifies the amount of ”mass” to be transported from $h_{i}$ to $\tilde{h_{i}}$. The function c($h_{i}$, $\tilde{h_{i}}$) is a cost function that quantifies the cost of transporting a unit of mass from $h_{i}$ to $\tilde{h_{i}}$. A common choice of c($h_{i}$, $\tilde{h_{j}}$) is the squared $l_{2}$ distance between the features, i.e., c($h_{i}$, $\tilde{h_{i}}$) = $||h_{i}-\tilde{h_{i}}||_{2}^{2}$, here, $1_{w\cdot h}$ is vector of ones. We also introduced entropy regularization [29] to reduce sensitivity to outlier instances.
Subsequently, the nearest neighbor feature for $H_{n}$ would be one $\tilde{H}_{i}$ in the CARLA dataset, which we define as $\tilde{H}^{*}$. Both $H_{n}$ and $\tilde{H}^{*}$ are then utilized for the subsequent process of feature merge. The formulation is as follows:
$\tilde{H}^{*}=\min\sum_{i=1}^{N_{s}}\operatorname{d_{OT}}\left(H_{n},\tilde{H}_{i}\right)$ (2)
@@ III-C Convex Merge and Fine-Tune
In the feature retrieval phase of autonomous driving, we have already obtained it through OT computed the optimal matching simulated feature $\tilde{H}^{*}$. Upon finding this match, we employ a convex merge operation to combine the real feature $H$ with the simulated feature $\tilde{H}^{*}$, resulting in a new composite feature $\hat{H}$, calculated as $\hat{H}=\pi(H,\tilde{H}^{*})$. The merging function $\pi$(·) is typically a convex combination method. The set of merged features denoted as $D_{m}=\{\hat{H_{i}}\}^{N_{m}}_{i=1}$, with $N_{m}$ being the number of features, is then utilized for the fine-tuning phase, aimed at enhancing the model’s ability to generalize from simulated to real data. The fine-tuning process is governed by the following loss function, which is consistent with the one used in Cross_view [6]:
$\mathcal{L}=\mathcal{L}_{BCE}+\lambda\mathcal{L}_{\text{cycle}}+\beta\left(\mathcal{L}_{1}^{D}+\mathcal{L}_{2}^{D}\right)$ (3)
During fine-tuning, we maintain the integrity of the pre-trained model by freezing all layers except for the decoder, which allows for the adaptation of the model to the new combined feature set without overwriting the learned representations.

### v2

@@ Abstract
As end-to-end autonomous driving advances toward real-world deployment, ensuring the safety of autonomous vehicles (AVs) has become a critical requirement for their commercial viability. While rule-based AVs have traditionally undergone rigorous testing in both real-world and simulated environments before deployment, data-driven autonomous models are typically trained on real-world datasets, limiting their generalization to simulation environments. This poses a significant challenge for the development and testing of end-to-end autonomous driving. To address this issue, we propose Retrieval-Augmented Learning for Autonomous Driving (RALAD), a novel framework designed to bridge the real-to-sim gap in a cost-effective manner. RALAD consists of three key components: (1) domain adaptation via an enhanced Optimal Transport (OT) method, which retrieves the most similar scenarios between real and simulated environments; (2) feature fusion across similar scenarios, enabling the construction of a feature mapping between real-world and simulated domains; and (3) feature extraction freezing with fine-tuning on the fused features, allowing the model to learn simulation-specific characteristics through feature mapping. We evaluate RALAD on three monocular 3D object detection models, and the results demonstrate that our approach significantly improves model accuracy in simulation. Additionally, we use real autonomous vehicle for testing in real-world scenarios, and have established simulated scenes similar to reality for further testing, which illustrate the effectiveness of our method.
@@ I INTRODUCTION
Figure 1: Performance of baseline model in simulated rainy weather: All models showed varying degrees of performance degradation under simulated rainy weather conditions.
@@
However, autonomous driving systems inevitably encounter corner cases, such as extreme weather conditions, unexpected pedestrian behavior, and rare road scenarios, which challenge their perception and decision-making capabilities [14, 15]. Given the high safety standards required in vehicle operation to prevent life-threatening accidents, it is imperative that autonomous systems rigorously address and test these corner cases to ensure robust performance [14, 15].
Models trained on real-world datasets predominantly encounter regular driving scenarios, making it challenging to cover and address the rare and complex corner cases [21, 16, 19]. To replicate and test corner cases, simulators such as CARLA and AirSim are commonly used to replicate and test specific scenarios. This approach has proven highly effective for rule-based autonomous driving in the past. [21]. However, with the rise of end-to-end autonomous driving, vision-based models impose significantly different granularity requirements on simulated and real-world scenes, resulting in a substantial domain gap. As shown in Figure 1, we collected a simulated dataset from CARLA and evaluated three models. All three models exhibited a performance decrease when tested on simulation, indicating that although these models perform well in the real-world, their performance tends to degrade in the simulated environment.
Current approaches to addressing this issue typically involve training the model directly on a mixture of real and simulated datasets [31] to mitigate the domain gap. However, these methods require full model retraining, incurring computational costs while potentially compromising the model’s original performance. Therefore, developing an efficient and minimally intrusive solution remains a significant challenge. While there are numerous domain generalization methods, many still require significant adjustments to the model. In contrast, retrieval-augmented learning (RAL) offers a more lightweight approach by enhancing the model’s ability to generalize across domains without the need for extensive retraining. In particular, methods using nearest neighbor search and feature fusion for domain adaptation [16, 18] inspire our approach to leveraging RAL to bridge the cross-domain gap between real and simulated datasets in autonomous driving.
While RAL has shown promise in domains like computer vision, applying this technique to autonomous driving introduces several challenges. Firstly, RAL was originally developed for classification tasks, whereas autonomous driving tasks often involve more complex, high-dimensional tasks like 3D detection, trajectory prediction, and decision-making, this creates a challenge in adapting RAL to these tasks. Furthermore, previous applications of RAL typically involved datasets from the same domain but different fields, such as Camelyon16 and Camelyon17. In contrast, real and simulated data in autonomous driving are fundamentally different in dimensions, which presents an additional challenge. Lastly, due to the sheer volume of autonomous driving data, efficiently adjust the model without affecting its original feature representations is a significant challenge.
We propose RALAD to address the gap between real and simulated images in autonomous driving. We introduce an enhanced Optimal Transport (OT) method for domain adaptation, which retrieves the most similar scenarios between real and simulated environments, and then through feature fusion, we construct a feature mapping between real and simulated feature maps. Subsequently, we utilize this feature mapping to allow the model to infer real-world features from simulated features. This approach eliminates the need for the model to learn features from simulated data, instead focusing on establishing the mapping between simulated and real features. Additionally, we freeze the feature extraction module, preserving the model’s original performance while significantly reducing the computational cost of training.
@@
  • We introduce RALAD, a framework utilizing OT method to establish a mapping between simulation and reality, in order to address the gap between reality and simulation in autonomous driving.
  • We apply RALAD to three models, achieving an improvement of up to 11.02% in accuracy on simulation while maintaining the accuracy on KITTI.
  • We use real autonomous vehicles and CARLA to create real and simulated environments, and conduct extensive experiments to validate the RALAD.
Figure 2: Overview of the RALAD framework for autonomous driving. The process begins with simulation and real-world datasets, which undergo Optimal Transport (OT) computation to retrieve the real feature map with the smallest transfer distance for the simulated feature map. The features from both domains are then mapped and merged, followed by fine-tuning on the combined features to enhance the model’s performance on autonomous driving (AD) tasks.
@@ II-A Retrieval-Augmented Learning
RAL is an approach that enhances learning models, it integrates retrieval mechanism to leverage existing data representations, thereby improving performance and efficiency. For example, Yottixel [22] employs a mix of supervised and unsupervised methods, including segmentation, clustering, and deep networks, to analyze image patches and employ distance metrics for retrieval. SISH [23] utilizes a tree structure for rapid WSI search and an uncertainty-based ranking for retrieval, reducing storage and labeling by building on preprocessed mosaics. HHOT [24] introduces optimal transport (OT) as a metric for comparing whole slide images (WSIs) or across WSI datasets, theoretically underpinning the application of OT for steering the retrieval and assembly of datasets. In RAM-MIL [18], the attention weight serves as a measure of probability density, signifying the ”mass” being transferred. By quantifying this, the method computes the conversion cost across various data domains. It then employs this distribution for nearest-neighbor retrieval, seamlessly integrating features from distinct domains to address out-of-domain challenges.
@@ II-B Gap Between Real And Sim
In the field of autonomous driving, there exists a gap between simulation and reality [25], which is caused by discrepancies in lighting, textures, vehicle dynamics, and agent behaviors. To address this, researchers have developed two primary approaches: knowledge transfer learning and digital twins (DTs) [17, 28]. In knowledge transfer learning, the RG problem is compounded by uneven environmental sampling and complex physical parameters [26]. To overcome this, researchers have developed strategies such as curriculum learning, meta-learning, and domain randomization [27]. Domain randomization, in particular, helps align simulation parameters with real-world variability, facilitating the transfer of learned strategies to real-world applications. Conversely, digital twin technology creates virtual models of real-world entities or systems. A case in point is the SynFog dataset [28], which uses an end-to-end process to produce photo-realistic synthetic data. Nonetheless, despite their promise, these methods confront the issue of high computational expenses, particularly in complex, dynamic real-world settings.
@@
3D object detection is crucial for autonomous driving as it enables vehicles to accurately perceive their surroundings, which is essential for safe navigation and decision-making. Conventionally, this has been achieved with the help of LiDAR sensors, which, although precise, are prohibitively expensive and computationally demanding. To address these limitations, the field has seen a significant advancement with the application of deep learning that leverage cameras for 3D detection. Specifically, the development of bird’s-eye view (BEV) representations from monocular images has emerged as a promising and more cost-effective alternative. Techniques such as MonoLayout [5] and CrossView [6] demonstrate the potential of using these BEV to perform 3D object detection. CrossView, in particular, has introduced a cross-view transformation module and a context-aware discriminator to enhance results, achieving cutting-edge performance in vehicle occupancy estimation. Nonetheless, issues such as class imbalance and low computational efficiency remain. The Dual-Cycled Cross-View Transformer network (DctNet) [7] has been proposed to tackle these challenges by integrating focal loss and optimizing multi-class learning, setting new standards of performance in the field of 3D object detection.
III METHODOLOGY
This section provides a detailed explanation of the RALAD framework, as shown in Figure 2. We introduce Retrieval-Augmented Learning based on Optimal Transport into autonomous driving.
@@ III-A Problem Formulation
The gap between real-world and simulation (real2sim) primarily arises from the cross-domain challenges between real and simulated data in autonomous driving. To address this, we consider two datasets: the Real dataset $D_{r}$ and the Sim dataset $D_{s}$. The Real dataset is defined as $D_{r}=\left\{X_{n},Y_{n}\right\}_{n=1}^{N_{r}}$, where $N_{r}$ is the number of images, and the Sim dataset is $D_{s}=\left\{\tilde{X_{m}},\tilde{Y_{m}}\right\}_{m=1}^{{N_{s}}}$, with ${N_{s}}$ being the number of simulated images, X and Y represent the image and ground truth, respectively. To bridge the gap between these two datasets, the first step is to establish a mapping relationship between them. We extract features from real and simulated images using an encoder function g(·), resulting in $H_{k}$ = g($X_{k}$) for real images and $\tilde{H}_{k}$ = g($\tilde{X}_{k}$) for simulated images. Our ultimate goal is to create pixel-to-pixel mapping relationship between features from different domain, thereby find the most similar real feature map from the real dataset for each simulated feature map. The formula is as follows:
$\mathcal{L}(\tilde{H}_{i},D_{r})=\min_{j=1}^{N_{r}}\operatorname{d_{OT}}\left(\tilde{H}_{i},H_{j}\right)$ (1)
III-B Retrieval-Augmented Learning based on Optimal Transport in Real2Sim
To address the mapping relationship between features from different domains, we introduce global feature level Optimal Transport, because the previous RAL method was mainly local and focused on the differences between individual instances in high-definition medical images, the instance images were relatively clear and did not have much interference. However, in autonomous driving, the environment has a significant impact on vehicle recognition, such as in rainy or snowy weather. So we proposed a global feature level Optimal Transport to globally quantify the differences between reality and simulation. As shown in Figure 3, we use the encoder of the pre-trained model to extract features from a single simulated image and all real images, and then separate the extracted features from each image. Then the simulated feature values are subjected to OT calculations with all real feature values. This ensures that a pixel-to-pixel mapping relationship can be established for simulating features to find more perfect real features, as shown in Figure 5, where vehicles and lanes almost correspond. This mapping process is repeated on the simulated dataset until the traversal of the simulated dataset is completed.
Figure 3: The one round calculation process of Optimal transport: the simulated image features of the single frame above and all the real image features below are separated, then each feature is multiplied by the same OT weights, and finally the simulated features of the single frame and all the real features are calculated for Optimal transport one by one.
@@ III-B Retrieval-Augmented Learning based on Optimal Transport in Real2Sim
During the OT calculation process, each image is treated as a probability distribution. We assign uniform weights to all instances and leave the non-uniform OT calculation to future exploration. The purpose of the OT algorithm is to calculate the distance between two features $H_{n}$ (real-world) and $\tilde{H_{m}}$ (simulation) for retrieval of the nearest feature. The formula is as follows:
$\operatorname{d_{OT}}\left(H_{n},\tilde{H_{m}}\right)=\sum_{i=1}^{w\cdot h}c\left(h_{ni},\tilde{h_{mi}}\right)T_{ii}+\beta\cdot\sum_{i}T_{ii}\log T_{ii}$ (2)
@@
In this equation, T denotes the transport plan matrix where each element $T_{ii}$ specifies the amount of ”mass” to be transported from $h_{ni}$ to $\tilde{h_{mi}}$, $h_{ni}$ and $\tilde{h_{mi}}$ are a feature value located at the real and simulated feature maps, respectively, w and h are the width and height of the feature map. The function c($h_{ni}$, $\tilde{h_{mi}}$) is a cost function that quantifies the cost of transporting a unit of mass from $h_{ni}$ to $\tilde{h_{mi}}$. A common choice of c($h_{ni}$, $\tilde{h_{mj}}$) is the squared $l_{2}$ distance between the features, i.e., c($h_{ni}$, $\tilde{h_{mi}}$) = $||h_{ni}-\tilde{h_{mi}}||_{2}^{2}$, here, $1_{w\cdot h}$ is vector of ones. We also introduced entropy regularization [29] to reduce sensitivity to outlier instances.
@@ III-C Convex Merge and Fine-Tune
In the feature retrieval phase of autonomous driving, we have already obtained it through OT computed the optimal matching simulated feature $\tilde{H}^{*}$. Upon finding this match, we employ a convex merge operation to combine the real feature $H$ with the simulated feature $\tilde{H}^{*}$, resulting in a new composite feature $\hat{H}$, calculated as $\hat{H}=\pi(H,\tilde{H}^{*})$, as shown in Figure 2. The merging function $\pi$(·) is typically a convex combination method, this fusion also allows us to establish mapping relationships on features. The set of merged features denoted as $D_{m}=\{\hat{H_{i}}\}^{N_{m}}_{i=1}$, with $N_{m}$ being the number of features, is then utilized for the fine-tuning phase, aimed at enhancing the model’s ability to generalize from simulated to real data.
During fine-tuning, we maintain the integrity of the pre-trained model by freezing all layers except for the decoder, because we previously used an encoder to extract features, then combined OT and Convex Merge to extract the mapping between real and simulated features, in order to maintain the ability to extract the mapping between real and simulated features, we froze the encoder layer of the model. Then, through this extracted mapping relationship, we only need to train the decoder layer of the model, so that the model can learn to establish a mapping relationship between reality and simulation, as shown in Figure 4.
Figure 4: The process of fine-tune: the model is frozen except for the decoder, and merged features are as inputs for the decoder to learn the mapping relationship between simulation and reality.
Figure 5: The table displays two sets of nearest neighbor results for KITTI 1, with these results drawn from the CARLA dataset. CARLA 1 stands out as the optimal match for KITTI 1 among the retrieved neighbors. And the table, a pair of bar graphs illustrate the transmission cost comparison between KITTI 1 and the CARLA features, a lower transmission cost indicates a higher degree of similarity between the images.